OpenStudy (anonymous):

The sum of two nonnegative numbers is 20. Find the numbers if the sum of their squares is as large as possible; as small as possible.

OpenStudy (anonymous):

x+y=20 A=x^2+y^2 A=x^2+(20-x)^2 A=2x^2-40x+400 A'=4x-40=0 x=10 y=10

OpenStudy (anonymous):

10^2+10^2=200 is min. 20^2+0^=400 max